OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (anonymous):

put it in y=mx+b form m=slope and b=the y intercept

OpenStudy (anonymous):

So what I normally do is draw out a graph of the line....

OpenStudy (anonymous):

basically what is meaning of slope in literaly

OpenStudy (anonymous):

but it says this y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

oohhh

OpenStudy (anonymous):

well first narrow down your answers that you know for sure are wrong

OpenStudy (anonymous):

exactly i dnt no which answer it is

OpenStudy (anonymous):

okay well hang on a second let me figure it out first then ill explain to you how to get the right aanswer... just a minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay so first, what I did was find the y=mx+b form of it from the info given, so do that and see if you get the same answer...

OpenStudy (anonymous):

ok i got that

OpenStudy (anonymous):

Did you get y=-6-15

OpenStudy (anonymous):

*-6x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now we have to take the possible answers and test to see which one is the correct equation... since the slope has to be -6 we can cancel out the 2nd and 4th answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

next we test both the equations by either simplifying them or testing points on the graph....